# MILR: Improving Multimodal Image Generation via Test-Time Latent Reasoning

### v2

@@ 4.2 Main Results
Figure 3: Qualitative studies on WISE. Reasoning cues are highlighted in red.
@@ 4.3.1 Hyperparameters
hu2025anchortokenmatchingimplicit have empirically shown that, in autoregressive image generation, early-stage tokens govern the overall image structure, and perturbing the first 20% of image tokens results in significant structural deviations. Therefore, we conservatively adopt a very small $\lambda_{v}$, varying it from 0.01 to 0.1 (see Figure 5 (b)). Surprisingly, optimizing only the first 2% of all image tokens already yields peak performance, while further increasing $\lambda_{v}$ tends to degrade it (see examples in Figure 12).
@@ 4.4 Reward Models
The reward model is a crucial component of MILR, providing learning signals for latent reasoning. Following previous work (liu2025flowgrpo; tong2025delving), we have used the benchmark’s evaluator as the reward model (denoted by [Uncaptioned image] OracleReward), but in real-world scenarios, oracle rewards are usually unknown, and it is difficult to design domain-specific rewards. To show that MILR is effective without the reliance on [Uncaptioned image] , we test it with a set of off-the-shelf reward models on GenEval:
@@
  • [Uncaptioned image] GPT-4o represents a frontier critic for assessing image quality (hurst2024gpt4o).
  • [Uncaptioned image] UnifiedReward is specifically tuned for a unified evaluation of MUG (wang2025unifiedreward).
  • [Uncaptioned image] MixedReward is a composite critic for more comprehensive evaluation. It aggregates rewards from specialized models, including GroundingDINO (liu2024grounding) (evaluating object detection), GIT (wang2022git) (judging colors), and [Uncaptioned image] (assessing aesthetics).
@@ 5 Discussion
We acknowledge two primary limitations of MILR. First, we have focused on an implementation of MILR that builds upon autoregressive MUG, where both text and image tokens are generated autoregressively. Another strong paradigm of MUG is through diffusion image generation, where image tokens within the same image can attend to each other (deng2025bagel). It is interesting to see if the effectiveness of MILR transfers. Second, MILR relies on a reward model for learning signals, but, in practice, a perfect reward model usually does not exist, and it is difficult to design a domain-agnostic reward model. In addition, our experimental results have revealed that the strongest non-oracle reward model still lags behind the oracle reward. Thus, future work is well-suited for designing reward models that can generalize like the unified reward model of wang2025unifiedreward.
@@ 6 Conclusion
Ethics Statement
We strictly adhere to the Code of Ethics. Below we elaborate on the ethical considerations relevant to our work and the measures we have put in place.
@@ A.1 Optimization algorithms
$\mathbf{z},\mathbf{t},\mathbf{v}\leftarrow p(\mathbf{t},\mathbf{v}|c)$ $\triangleright$ Initial latent vectors: Equation 1
@@
Note. $\mathcal{J}(\mathbf{z})$ as defined in  Equation 6; $V_{f}$ sampled as in  Equation 4; $T_{f}$ is the full reasoning text. $\mathbf{z}^{(t)}_{k}$ and $\mathbf{z}^{(v)}_{k}$ denote the text and visual latents at iteration $k$.
@@ A.3.2 Analysis of Reward models
In Section 4.4, we evaluated MILR with different reward models: [Uncaptioned image] SelfReward, [Uncaptioned image] GPT-4o, [Uncaptioned image] UnifiedReward, [Uncaptioned image] MixedReward, and [Uncaptioned image] OracleReward. Table 6, Table 7 and Table 8 show results across benchmarks with different rewards. We further provide example images generated under each setup in Figure 13. Overall, [Uncaptioned image] establishes an upper bar of generation quality, and [Uncaptioned image] , among non-oracle reward models, performs best because it incorporates multiple specialized critics to provide a comprehensive assessment. Though [Uncaptioned image] still lags behind [Uncaptioned image] , in the absence of [Uncaptioned image] , we can compose a strong universal reward model from specialized off-the-shelf image critics.
@@
Figure 13: Case study of reward models.
@@ A.3.3 Efficieny Analysis
We summarize the computing resources used by different image generation models in Table 9. Compared with training-based reasoning approaches such as T2I-R1 (jiang2025t2i-r1), Janus-Pro with DPO/GRPO (tong2025delving), and Flow-GRPO (liu2025flowgrpo), our test-time reasoning method achieves the best performance, without relying on curated reasoning data for training. As an example, Flow-GRPO requires about 2K A800 GPU hours for training but only matches MILR, not to mention its labor cost in curating reasoning data. Moreover, when compared with the strong test-time reasoning method Best-of-N (N=20 vs. T=20 in MILR), MILR surpasses it by 0.04 while requiring less inference time because it employs an early-stop strategy, i.e., it stops once the generated image satisfies evaluation criteria.
@@
T2I-R1 (jiang2025t2i-r1) 8 H800 16 h 0.79
Janus-Pro+GRPO (tong2025delving) 8 A100 $\sim$9 h 0.81
Janus-Pro+DPO (tong2025delving) 8 A100 $\sim$9 h 0.83
Flow-GRPO (liu2025flowgrpo) 24 A800 $\sim$100 h 0.95
Reflect-DiT (li2025reflectdit) A6000 24 h 16 h 0.81
@@
A.4 The Use of Large Language Models (LLMs)

### v3

@@ 4.2 Main Results
Figure 3: Qualitative studies on theree benchmarks. Reasoning cues are highlighted in red.
@@ 4.3.1 Hyperparameters
Hu et al. (2025) have empirically shown that, in autoregressive image generation, early-stage tokens govern the overall image structure, and perturbing the first 20% of image tokens results in significant structural deviations. Therefore, we conservatively adopt a very small $\lambda_{v}$, varying it from 0.01 to 0.1 (see Figure 5 (b)). Surprisingly, optimizing only the first 2% of all image tokens already yields peak performance, while further increasing $\lambda_{v}$ tends to degrade it (see examples in Figure 12).
@@ 4.4 Reward Models
The reward model is a crucial component of MILR, providing learning signals for latent reasoning. Following previous work (Liu et al., 2025a; Tong et al., 2025), we have used the benchmark’s evaluator as the reward model (denoted by [Uncaptioned image] OracleReward), but in real-world scenarios, oracle rewards are usually unknown, and it is difficult to design domain-specific rewards. To show that MILR is effective without the reliance on [Uncaptioned image] , we test it with a set of off-the-shelf reward models on GenEval:
@@
  • [Uncaptioned image] GPT-4o represents a frontier critic for assessing image quality (Hurst et al., 2024).
  • [Uncaptioned image] UnifiedReward is specifically tuned for a unified evaluation of MUG (Wang et al., 2025).
  • [Uncaptioned image] MixedReward is a composite critic for more comprehensive evaluation. It aggregates rewards from specialized models, including GroundingDINO (Liu et al., 2024) (evaluating object detection), GIT (Wang et al., 2022) (judging colors), and [Uncaptioned image] (assessing aesthetics).
@@ 5 Discussion
We acknowledge two primary limitations of MILR. First, we have focused on an implementation of MILR that builds upon autoregressive MUG, where both text and image tokens are generated autoregressively. Another strong paradigm of MUG is through diffusion image generation, where image tokens within the same image can attend to each other (Deng et al., 2025). It is interesting to see if the effectiveness of MILR transfers. Second, MILR relies on a reward model for learning signals, but, in practice, a perfect reward model usually does not exist, and it is difficult to design a domain-agnostic reward model. In addition, our experimental results have revealed that the strongest non-oracle reward model still lags behind the oracle reward. Thus, future work is well-suited for designing reward models that can generalize like the unified reward model of Wang et al. (2025).
@@ 6 Conclusion
7 Ethics Statement
We strictly adhere to the ICLR Code of Ethics. Below we elaborate on the ethical considerations relevant to our work and the measures we have put in place.
@@ 7 Ethics Statement
8 Reproducibility statement
To ensure reproducibility, we present experimental details, including settings and hyperparameter configurations, in Section 4.1 and release our code at https://github.com/spatigen/milr.
@@ A.1 Optimization algorithms
$\mathbf{z},\mathbf{t},\mathbf{v}\leftarrow p(\mathbf{t},\mathbf{v}|c)$ $\triangleright$ Initial latent vectors
@@ A.3.2 Analysis of Reward models
In Section 4.4, we evaluated MILR with different reward models: [Uncaptioned image] SelfReward, [Uncaptioned image] GPT-4o, [Uncaptioned image] UnifiedReward, [Uncaptioned image] MixedReward, and [Uncaptioned image] OracleReward (see Table 6, Table 7 and Table 8). We further provide example images generated under each setup in Figure 13. Overall, [Uncaptioned image] establishes an upper bar of generation quality, and [Uncaptioned image] , among non-oracle reward models, performs best because it incorporates multiple specialized critics to provide a comprehensive assessment. Though [Uncaptioned image] still lags behind [Uncaptioned image] , in the absence of [Uncaptioned image] , we can compose a strong universal reward model from specialized off-the-shelf image critics.
@@
Figure 13: Case study of various reward models.
@@ A.3.3 Efficieny Analysis
We summarize the computing resources used by different image generation models in Table 9. Compared with training-based reasoning approaches such as T2I-R1 (Jiang et al., 2025a), Janus-Pro with DPO/GRPO (Tong et al., 2025), and Flow-GRPO (Liu et al., 2025a), our test-time reasoning method achieves the best performance, without relying on curated reasoning data for training. As an example, Flow-GRPO requires about 2K A800 GPU hours for training but only matches MILR, not to mention its labor cost in curating reasoning data. Moreover, when compared with the strong test-time reasoning method Best-of-N (N=20 vs. T=20 in MILR), MILR surpasses it by 0.04 while requiring less inference time because it employs an early-stop strategy, i.e., it stops once the generated image satisfies evaluation criteria.
@@
T2I-R1 (Jiang et al., 2025a) 8 H800 16 h 0.79
Janus-Pro+GRPO (Tong et al., 2025) 8 A100 $\sim$9 h 0.81
Janus-Pro+DPO (Tong et al., 2025) 8 A100 $\sim$9 h 0.83
Flow-GRPO (Liu et al., 2025a) 24 A800 $\sim$100 h 0.95
Reflect-DiT (Li et al., 2025b) A6000 24 h 16 h 0.81
@@
A.4 Latent Regularization to Mitigate Reward Hacking
As discussed in Section 4.4.1, MILR can suffer from reward hacking, which refers to a scenario where a MUG model achieves high rewards but exhibits low image generation quality. To mitigate this issue, we adopt a simple $\ell_{2}$ regularizer to constrain the difference between the latents before and after optimization. Let $\mathbf{z}_{\text{init}}$ denote the initial latents produced by a frozen MUG model before optimization, and let $\mathcal{J}(\mathbf{z}^{k})$ denote the original objective used at the latents optimization step $k$ (see Equation 6), we define the regularized objective as:
$\mathcal{J}_{\text{reg}}(\mathbf{z}^{k})=\mathcal{J}(\mathbf{z}^{k})-\beta\big\|\mathbf{z}^{k}-\mathbf{z}_{\text{init}}\big\|_{2}^{2}\,$, (7)
where $\beta>0$ controls the strength of regularization. Intuitively, this term encourages the optimized latents $\mathbf{z}^{k}$ to stay close to the initial latents, thus maintaining image generation capability similar to that of the frozen MUG model.
In our experiments, we use a small coefficient $\beta=0.05$, which empirically stabilizes the optimization and alleviates semantic collapse, without degrading generation quality. In Figure 14, we illustrate that adding latent regularization helps mitigate reward hacking.
Figure 14: $\ell_{2}$ regularization on latents mitigates reward hacking.
A.5 The Use of Large Language Models (LLMs)